# Do quantum linear solvers offer advantage for networks-based system of linear equations?

### v3

@@ Abstract
In this exploratory numerical study, we assess the suitability of Quantum Linear Solvers (QLSs) toward providing a quantum advantage for Networks-based Linear System Problems (NLSPs). NLSPs naturally arise from graphs, and are of importance as they are connected to real-world applications. The achievable advantage with a QLS for an NLSP depends on the interplay between the scaling of condition number and sparsity of matrices associated with the graph family considered, as well as system size growth. We analyze 50 graph families and identify that within the scope of our study, only 4% of them exhibit prospects for an exponential advantage with the Harrow-Hassidim-Lloyd (HHL) algorithm relative to an efficient classical solver (best graphs), while about 20% of them show a polynomial advantage (better graphs). Furthermore, we report that some graph families graduate from offering no advantage with HHL to promising a polynomial advantage with improved algorithms such as the Childs-Kothari-Somma algorithm, while some other graph families exhibit futile exponential advantage. We introduce a unified graph superfamily and show the existence of infinite best and better graphs in it. We also conjecture the conditions under which one may visually examine a graph family and guess the prospects for an advantage. Finally, we very briefly touch upon some practical issues that may arise even if the aforementioned graph theoretic requirements are satisfied, including quantum hardware challenges.
@@ I Introduction
Quantum algorithms promise speed-up for certain problems relative to their best known classical counterparts, thus motivating the ongoing second quantum revolution, which involves building reliable quantum computers to eventually advance toward commercial realization [scholten2024assessing, bova2021commercial, gill2025quantum, macquarrie2020emerging, cumming2022using, sinno2023performance, ruane2025quantum]. Quantum linear solvers (QLSs) are particularly significant in this context, as they are among the few classes of known quantum algorithms that can, in principle, offer an exponential advantage [morales]. A QLS prescribes a protocol towards evaluating systems of linear equations, $A\vec{x}=\vec{b}$, where the $(\mathcal{N}\times\mathcal{N})$ matrix $A$ and the vector $\vec{b}$ are known quantities. Since such systems are ubiquitous in natural sciences and engineering [anton2013elementary, machol1961linear, hamid2019balancing, wilson1958solution] solving them efficiently in the context of leveraging the aforementioned quantum speed-up is of importance. An example of a QLS is the well-known Harrow-Hassidim-Lloyd (HHL) algorithm [Harrow2009QuantumEquations], whose runtime complexity goes as $poly(\mathrm{log}(\mathcal{N}),s,\kappa,1/\epsilon)$ where $\mathcal{N}$, $s$, and $\kappa$ refer to the system size, sparsity, and the condition number of the $A$ matrix respectively, while $\epsilon$ refers to the additive error in the output state after performing HHL. The algorithm can outperform their best known classical counterparts such as the conjugate gradient algorithm [cg1994], when these parameters grow in a certain way with respect to the system size. However, identifying such settings where QLSs offer such an advantage is challenging in spite of a lot of efforts in literature in this direction. Our exploratory work is aimed toward addressing this timely problem.
@@
Analysis of quantum advantage from HHL for specific problems– Prior numerical studies in this direction indicate that the prospects of speed-up is limited at best due to the condition number scaling. In Ref. [coffrin], the authors consider the DC power flow problem and analyze the end-to-end complexity of the HHL algorithm, including obtaining the scaling behaviours of the condition number and sparsity with system size for this application. Their numerical simulations demonstrate that since $\kappa$ grows polynomially in system size (where system size is the number of buses), practical advantage from HHL for this application is unlikely. The authors of Ref. [mlhhl] study the suitability of the HHL algorithm along these lines in the context of labeling problems using machine learning classifiers, and reach the conclusion that the condition number has a critical impact on the problem. The authors of Ref. [hydro] explore modeling hydrological fracture networks, and en route, carry out an analysis of $\kappa$ with system size. Their results point to $\kappa$ growing unfavourably in system size, unless preconditioning, which itself is classically resource intensive, is employed. A work by the authors of Ref. [yalovetzky2021hybrid] applies the HHL algorithm to finance (portfolio optimization), and their data indicates that $\kappa$ scales quadratically with system size(number of assets) even in the best case scenario. A relatively recent work that comments on $\kappa$ scaling is Ref. [gopalakrishnan2024solving], where the authors solve Hele-Shaw flow in fluid dynamics using HHL, and find that $\kappa$ scales exponentially for their example case with system size (domain grid points).
@@
Analysis of quantum advantage from HHL in the most general setting– The exact opposite viewpoint would involve comparing the runtime complexities of the HHL and some suitably chosen efficient classical algorithm based on their respective complexity expressions. The authors of Ref. [tu2025towards] do exactly this in their work, with the end goal of carrying out a resource estimation analysis in terms of space, time, and energy for the HHL algorithm.
@@
Figure 1: An overview of the current study. (a) Illustration of the linear systems problem (where for simplicity, we assume real-valued entries for $A$, $\vec{b}$, and $\vec{x}$) and the runtime complexity scaling of quantum linear solvers and an efficient classical linear solver, for which we happen to borrow the runtime expression of the otherwise limited conjugate gradient method. Here, ${f}(s(\mathcal{N}))$ denotes a function of sparsity, $s(\mathcal{N})$, which in turn depends on the system size, $\mathcal{N}$. (b) Depiction of the connection between real-world applications, such as effective resistance determination and traffic flow congestion detection with graph Laplacian and graph incidence matrices respectively, and linear equations. (c) Schematic of our numerical survey on 50 graph families, where for each of them, we study $\kappa(\mathcal{N})$ and $s(\mathcal{N})$ behaviour with system size, $\mathcal{N}$ (NN for Laplacian matrix and $N+M$, for the incidence matrix) as well as system size scaling with $n$, where $n=1,2,\cdots$, to infer within the scope of our calculations if there is a potential for exponential advantage (best graph family), polynomial advantage (better graph family), sub-linear advantage (good graph family), or no advantage (bad graph family), all with the HHL algorithm and compared relative to the efficient classical linear solver.
@@
Aurea mediocritas?: Analysis of quantum advantage from quantum linear solvers in networks-based linear system problems– A third route, a road not taken, could be to chart a middle course by neither pursuing application-specific studies nor a fully general analysis, but rather group together diverse applications under a common mathematical framework, and then perform extensive numerical analysis to analyze the potential for advantage. Such a framework should offer flexible scaling in $\kappa$ and $s$, and system size. While HHL-based quantum chemistry application [psihhl] seems promising, showing polylogarithmic scaling in $\kappa$ with system size (number of particle-hole excitations arising from a reference state) for limited set of molecules, the structure of the underlying quantum many-body theoretic equations constrains the $A$ matrix growth to a polynomial function in system size, and thus the framework is not sufficiently flexible for our considerations.
@@
For this work, we consider for our numerical analyses the highly flexible Networks-based Linear System Problems (NLSP) framework, where one begins with a complex network and by applying a set of rules arrives at a system of linear equations. The NLSP framework accommodates a panoply of graphs/complex networks, and thus admits not only a huge variety of functions for $\kappa$, $s$, and system size, but also many potential applications. We pick two types of NLSPs, one where the $A$ matrix is the graph Laplacian, and the other where it is the graph incidence matrix. The former finds its applications in problems such as determining effective resistances in electrical circuits and finding voltages in power flow problems [vishnoi2013lx, spielman2010algorithms], while the latter can be used in applications that involve finding flows in branches of networks, such as traffic flow congestion detection [anton2013elementary].
@@
Our study investigates different graph families to identify those that offer an advantage. Through detailed numerical analyses, we study the scaling of $\kappa$, which has garnered much attention in literature as discussed in the earlier paragraphs, $s$, and system size growth across candidate graph families. Using this data, we compare the runtime of a QLS relative to a fictitious efficient classical linear solver (CLS), enabling us to estimate the crossover points where quantum advantage can be realized.
Following our identification of best, better, good and bad graph families from our numerical analyses, we take the route of building larger graph families from existing ones; we do so via the generalized hypercube graph family, where we identify an infinity of best and better graph families.
@@
As performing numerical analyses to categorize the graph families are computationally costly, we ask whether one can qualitatively assess $\kappa$, $s$, and system size growth and thus by extension the possibility of an advantage by only looking at a graph family. We conjecture that it is possible, based on conclusions from our data.
@@
The rest of the work is structured as follows: Given that the topics discussed in this study lie at the intersection of quantum algorithms, complexity, and graph theory and thus can be of interest to readers from all three communities, we attempt to make the article self-contained by devoting Sections II and III for introducing quantum linear solvers and network-based linear system problems. Our results and subsequent discussions form the remaining sections: In Section IV, we discuss the results from our survey on candidate complex networks and their suitability for achieving quantum advantage. We begin with the results for Laplacian matrix (Section IV.1), and then move to incidence matrix (Section IV.2). This is followed by section V, where we discuss a general graph superfamily construction from which we identify new best and better graphs. We discuss our conjecture to guess the possibility of advantage from a graph construction in VI. In Section VII, we briefly comment on our quantum hardware computations carried out on toy matrices. Finally, we conclude with a summary of the work and future prospects in Section VIII.
@@ II.1 Definitions
In this work, we consider the following functions for system size growth:
@@ II.2 The HHL algorithm
As discussed earlier, $\mathcal{N}$ is the system size, $s(\mathcal{N})$ is the sparsity of $A$, $\kappa(\mathcal{N})$ the condition number of $A$, and $\epsilon(\mathcal{N})$ is the error in the output state that we incur in the algorithm. In deriving the complexity of HHL, this error is assumed to be solely from inadequacy in the number of clock register qubits [Harrow2009QuantumEquations]. We note that hereafter, whenever we mention $\mathrm{log}$, it is the natural logarithm. We also note that we have explicitly shown the dependence of condition number, sparsity, and error on $\mathcal{N}$ to stress its importance in the context of our study. The scaling in $\kappa(\mathcal{N})$ and $\epsilon(\mathcal{N})$ are usually considered as drawbacks, and variants of HHL and subsequent QLSs improve on one or both of these aspects.
@@ II.3 Variants of HHL and other QLSs
We now list three variants of HHL that we consider for this work, all of which focus on reducing $\kappa$ scaling:
  • HHL with Amplitude Amplification (HHL-AA) [Harrow2009QuantumEquations, Brassard]: This is often assumed when one discusses HHL, but since the subroutine adds significant depth to the HHL quantum circuit (for example, see Ref. [morales]), we keep it distinct from the original HHL circuit. The benefit that the variant offers is a reduction of complexity in $\kappa(\mathcal{N})$, from $\kappa(\mathcal{N})^{3}$ to $\kappa(\mathcal{N})^{2}$. Practically, this has the effect of reducing the number of shots in an HHL calculation.
  • HHL with Variable Time Amplitude Amplification (HHL-VTAA) [ambainis]: This approach improves over HHL-AA and can be thought of as its generalization. The method reduces the complexity further to $\kappa(\mathcal{N})\mathrm{log}^{3}(\kappa(\mathcal{N}))$, but trades off in $\epsilon(\mathcal{N})$ scaling.
  • Psi-HHL [psihhl]: This approach reduces the complexity in $\kappa$ to its optimal scaling, that is, $\kappa(\mathcal{N})$, and with little increase in circuit depth.
We also list the other QLSs that we consider besides HHL and its variants:
  • Phase randomisation method [prm]: The method is inspired by the adiabatic quantum computing model, and employs evolution randomisation. The method scales as
@@
  • AQC(exp) method [toa]: The work demonstrates solving system of linear equations within the adiabatic quantum computing framework. This method too scales as
@@
    and thus is near-optimal in scaling of both condition number and precision, but unlike the CKS algorithm, the use of the expensive VTAA step is avoided.
@@
    The dream QLS serves as a benchmark to how much of an advantage we can get in the best case scenario, and thus subsumes all other QLSs that we do not consider. We assume in defining this solver that one cannot go below $\kappa(\mathcal{N})$ [Harrow2009QuantumEquations], $\sqrt{s(\mathcal{N})}$, and $\mathrm{log}(1/\epsilon(\mathcal{N}))$ [morales] in its complexity expression.
@@
HHL [Harrow2009QuantumEquations] $\mathcal{O}\textbf{\bigg(}\mathrm{log}(\mathcal{N})s(\mathcal{N})^{2}\kappa(\mathcal{N})^{3}\frac{1}{\epsilon(\mathcal{N})}\textbf{\bigg)}$
HHL-AA [Harrow2009QuantumEquations, Brassard] $\mathcal{O}\textbf{\bigg(}\mathrm{log}(\mathcal{N})s(\mathcal{N})^{2}\kappa(\mathcal{N})^{2}\frac{1}{\epsilon(\mathcal{N})}\textbf{\bigg)}$
HHL-VTAA [ambainis] $\mathcal{O}\textbf{\bigg(}\mathrm{log}(\mathcal{N})s(\mathcal{N})^{2}\kappa(\mathcal{N})\mathrm{log}^{3}\bigg(\frac{\kappa(\mathcal{N})}{\epsilon(\mathcal{N})}\bigg)\frac{1}{\epsilon(\mathcal{N})^{3}}\mathrm{log}^{2}\bigg(\frac{1}{\epsilon(\mathcal{N})}\bigg)\textbf{\bigg)}$
Psi-HHL [psihhl] $\mathcal{O}\textbf{\bigg(}\mathrm{log}(\mathcal{N})s(\mathcal{N})^{2}\kappa(\mathcal{N})\frac{1}{\epsilon(\mathcal{N})}\textbf{\bigg)}$
Phase randomisation method [prm] $\mathcal{O}\textbf{\bigg(}\mathrm{log}(\mathcal{N})s(\mathcal{N})\kappa(\mathcal{N})\mathrm{log}(\kappa(\mathcal{N}))\frac{1}{\epsilon(\mathcal{N})}\textbf{\bigg)}$
CKS algorithm [Childs] $\mathcal{O}\textbf{\bigg(}\mathrm{log}(\mathcal{N})s(\mathcal{N})\kappa(\mathcal{N})\mathrm{polylog}\bigg(s(\mathcal{N})\kappa(\mathcal{N})\frac{1}{\epsilon(\mathcal{N})}\bigg)\textbf{\bigg)}$
AQC(exp) method [toa] $\mathcal{O}\textbf{\bigg(}\mathrm{log}(\mathcal{N})s(\mathcal{N})\kappa(\mathcal{N})\mathrm{polylog}\bigg(s(\mathcal{N})\kappa(\mathcal{N})\frac{1}{\epsilon(\mathcal{N})}\bigg)\textbf{\bigg)}$
@@
Table 1 presents the expression for the runtime complexity of each of the aforementioned quantum algorithms in its second column. We recall that we assume in the table and also throughout hereafter that $\epsilon^{-1}\sim\mathrm{log}(\mathcal{N})$. We also drop the $\mathcal{O}$ hereafter when we discuss complexity expressions, for brevity.
@@
The specific functional forms for ${f}(s(\mathcal{N}))$ and ${g}(\kappa(\mathcal{N}))$ depend on the QLS chosen. On the other hand, $\kappa(\mathcal{N})$, $s(\mathcal{N})$, and $\mathcal{N}(n)$ depend on the choice of application, and they determine the degree of quantum advantage that one obtains for an application. The parameter $n\in\mathbb{N}$ defines the way system size grows, that is, $\mathcal{N}=\mathcal{N}(n)$. To that end, we define another useful quantity:
Definition II.5.
$\tilde{R}(n)=\frac{t_{\mathrm{CLS}}(n)}{t_{\mathrm{QLS}}(n)}.$
We make some general observations at this point based on Definition II.4. We shall revisit Definition II.5 in Section IV. When $R(\mathcal{N})>1$, a QLS under consideration performs better than CLS. That is, when
@@
That is, if an $(\mathcal{N}\times\mathcal{N})$ matrix $A$ is ill-conditioned enough and/or is sufficiently dense such that the product $\kappa(\mathcal{N})^{5/2}s(\mathcal{N})$ is equal to or greater than the right hand side, then we cannot expect an advantage from HHL. If we go to an improved variant such as Psi-HHL, the product $\kappa(\mathcal{N})^{1/2}s(\mathcal{N})$ can be less than the right hand side for a matrix $A$ that is relatively denser and/or relatively more ill-conditioned. The condition for the phase randomisation method is particularly interesting, as its sparsity scaling is the same as that of CLS, and thus it is only the condition number scaling that matters when we compare the two approaches. We also note that for the cases of the CKS and the AQC(exp) algorithms, it is not possible to separate out $\kappa(\mathcal{N})$ and $s(\mathcal{N})$ to one side due to the functional form for their complexities. Lastly, it is worth adding that since the complexity expressions contain a polylog in them, we will, for the purposes of our numerical analyses in the subsequent sections, consider three cases for the CKS and AQC(exp) algorithms for polylog: log, log2, and log3 behaviours. In the subsequent figures as well as main text, we use a shorthand notation: for example, the AQC(exp) algorithm with the polylog chosen to be log is denoted as AQC(1), and so on.

### v4

@@ Abstract
In this exploratory numerical study, we assess the suitability of Quantum Linear Solvers (QLSs) toward providing a quantum advantage for Networks-based Linear System Problems (NLSPs). NLSPs naturally arise from graphs, and are of importance as they are connected to real-world applications. The achievable advantage with a QLS for an NLSP depends on the interplay between the scaling of condition number and sparsity of matrices associated with the graph family. We analyze 50 graph families and identify that within the scope of our study, only 21 of them exhibit prospects for an exponential advantage with the Harrow-Hassidim-Lloyd (HHL) algorithm relative to an efficient classical solver. We call graph families that offer advantage with HHL as good graph families. We also compare the performance of the considered 50 graph families with 7 other QLSs. Furthermore, we report that some graph families graduate from offering no advantage with HHL to promising an exponential advantage with improved algorithms such as the Childs-Kothari-Somma algorithm. We also introduce a unified graph superfamily and show the existence of infinite good graph families in it. Since the runtime expressions for linear solvers involve condition number, which in itself is not easy to compute, ascertaining advantage prospects with quantum linear solvers itself is not an easy problem. Thus, we conjecture the conditions under which one may visually examine a graph family and guess the prospects for an advantage. Finally, we very briefly touch upon some practical issues that may arise even if the aforementioned graph theoretic requirements are satisfied, including quantum hardware challenges.
@@ I Introduction
Quantum algorithms promise speed-up for certain problems relative to their best known classical counterparts, thus motivating the ongoing second quantum revolution, which involves building reliable quantum computers to eventually advance toward commercial realization [56, 9, 25, 43, 18, 58, 55]. Quantum linear solvers (QLSs) are particularly significant in this context, as they are among the few classes of known quantum algorithms that can, in principle, offer an exponential advantage [46]. A QLS prescribes a protocol towards evaluating systems of linear equations, $A\vec{x}=\vec{b}$, where the $(\mathcal{N}\times\mathcal{N})$ matrix $A$ and the vector $\vec{b}$ are known quantities. Since such systems are ubiquitous in natural sciences and engineering [4, 42, 29, 67], solving them efficiently in view of the aforementioned quantum speed-up is of importance. An example of a QLS is the well-known Harrow-Hassidim-Lloyd (HHL) algorithm [31], whose runtime complexity goes as $\mathrm{poly}(\mathrm{log}(\mathcal{N}),s,\kappa,1/\epsilon)$ where $\mathcal{N}$, $s$, and $\kappa$ refer to the system size, sparsity, and the condition number of the $A$ matrix respectively, while $\epsilon$ refers to the additive error in the output state after performing HHL. The algorithm can outperform their best known classical counterparts such as the conjugate gradient algorithm [57], when these parameters grow in a certain way with respect to the system size. However, identifying such settings where QLSs offer such an advantage is challenging in spite of a lot of efforts in literature in this direction. Our exploratory work is aimed toward addressing this timely problem.
@@
Analysis of quantum advantage from HHL for specific problems– Prior numerical studies in this direction indicate that the prospects of speed-up are limited at best due to the condition number scaling. In Ref. [50], the authors consider the DC power flow problem and analyze the end-to-end complexity of the HHL algorithm, including obtaining the scaling behaviours of the condition number and sparsity with system size for this application. Their numerical simulations demonstrate that since $\kappa$ grows polynomially in system size (where system size is the number of buses), practical advantage from HHL for this application is unlikely. The authors of Ref. [19] study the suitability of the HHL algorithm along these lines in the context of labeling problems using machine learning classifiers, and reach the conclusion that the condition number has a critical impact on the problem. The authors of Ref. [26] explore modeling hydrological fracture networks, and en route, carry out an analysis of $\kappa$ with system size. Their results point to $\kappa$ growing unfavourably in system size, unless preconditioning, which itself is classically resource intensive, is employed. A work by the authors of Ref. [68] applies the HHL algorithm to finance (portfolio optimization), and their data indicates that $\kappa$ scales quadratically with system size(number of assets) even in the best case scenario. A relatively recent work that comments on $\kappa$ scaling is Ref. [27], where the authors solve Hele-Shaw flow in fluid dynamics using HHL, and find that $\kappa$ scales exponentially for their example case with system size (domain grid points). A recent analysis on limited molecular systems indicated that HHL-based quantum chemistry application [64] seems promising as it shows polylogarithmic scaling in $\kappa$.
@@
Analysis of quantum advantage from HHL in the most general setting– The exact opposite viewpoint would involve comparing the runtime complexities of the HHL and some suitably chosen efficient classical algorithm based on their respective complexity expressions. The authors of Ref. [65] do exactly this in their work, with the end goal of carrying out a resource estimation analysis in terms of space, time, and energy for the HHL algorithm.
@@
Figure 1: An overview of the current study. (a) Illustration of the linear systems problem (where for simplicity, we assume real-valued entries for $A$, $\vec{b}$, and $\vec{x}$) and the runtime complexity scaling of quantum linear solvers and an efficient classical linear solver, for which we happen to borrow the runtime expression of the otherwise limited conjugate gradient method. Here, ${f}(s(\mathcal{N}))$ denotes a function of sparsity, $s(\mathcal{N})$, which in turn depends on the system size, $\mathcal{N}$. (b) Depiction of the connection between real-world applications, such as effective resistance determination and traffic flow congestion detection with graph Laplacian and graph incidence matrices respectively, and linear equations. (c) Schematic of our numerical survey on 50 graph families, where for each of them, we study $\kappa(\mathcal{N})$ and $s(\mathcal{N})$ behaviour with system size, $\mathcal{N}$ (NN for Laplacian matrix and $N+M$, for the incidence matrix), to infer within the scope of our calculations the potential for quantum advantage (good graph family) or no advantage (bad graph family), all with the HHL algorithm and compared relative to the efficient classical linear solver.
@@
Aurea mediocritas?: Analysis of quantum advantage from quantum linear solvers in networks-based linear system problems– A third route, a road not taken, could be to chart a middle course by neither pursuing application-specific studies nor a fully general analysis, but rather group together diverse applications under a common mathematical framework, and then perform extensive numerical analysis to analyze the potential for advantage. Such a framework should offer flexible scaling in $\kappa$ and $s$, and system size.
@@
For this work, we consider for our numerical analyses the highly flexible Networks-based Linear System Problems (NLSP) framework, where one begins with a complex network and by applying a set of rules arrives at a system of linear equations. The NLSP framework accommodates a panoply of graphs/complex networks, and thus admits not only a huge variety of functions for $\kappa$, $s$, and system size, but also many potential applications. We pick two types of NLSPs, one where the $A$ matrix is the graph Laplacian, and the other where it is the graph incidence matrix. The former finds its applications in problems such as determining effective resistances in electrical circuits and finding voltages in power flow problems [66, 60], while the latter can be used in applications that involve finding flows in branches of networks, such as traffic flow congestion detection [4].
@@
Our study investigates different graph families to identify those that offer an advantage. Through detailed numerical analyses, we study the scaling of $\kappa$, which has garnered much attention in literature as discussed in the earlier paragraphs, and $s$ across candidate graph families. It is especially important to note that finding analytical expressions for $\kappa$ for a graph family is extremely rare, and is hard in general. This necessitates and motivates our numerical study. Using our numerical data for $\kappa$ and $s$, we compare the runtime of a QLS relative to a fictitious efficient classical linear solver (CLS), enabling us to estimate the crossover points where quantum advantage can be realized. We identify graph families that offer advantage with HHL (good graph families), and also discuss a new graph superfamily that subsumes in it an infinity of good graph families.
@@
As performing numerical analyses to categorize the graph families are computationally costly, we ask whether one can qualitatively assess $\kappa$ and $s$, and thus by extension the possibility of assessing prospects of an advantage by only looking at small instances of a graph family. We conjecture that it is possible, based on conclusions from our data.
@@
The rest of the work is structured as follows: Given that the topics discussed in this study lie at the intersection of quantum algorithms, complexity, and graph theory and thus can be of interest to readers from all three communities, we attempt to make the article self-contained by devoting Sections II and III for introducing quantum linear solvers and networks-based linear system problems. Our results and subsequent discussions form the remaining sections: In Section IV, we discuss the results from our survey on candidate complex networks and their suitability for achieving quantum advantage. We begin with the results for Laplacian matrix (Section IV.1), and then move to incidence matrix (Section IV.2). This is followed by Section V, where we discuss a general graph superfamily construction from which we identify new good graphs. We introduce a conjecture to guess the possibility of advantage from a graph construction in Section VI. In Section VII, we briefly comment on our quantum hardware computations carried out on toy matrices. Finally, we conclude with a summary of the work and future prospects in Section VIII.
@@ II.1 Definitions
In this work, we consider the following functions for growth of $\kappa$ and $s$ with system size:
@@ II.2 The HHL algorithm
As discussed earlier, $\mathcal{N}$ is the system size, $s(\mathcal{N})$ is the sparsity of $A$, $\kappa(\mathcal{N})$ the condition number of $A$, and $\epsilon(\mathcal{N})$ is the additive error in the output state that we incur in the algorithm. In deriving the complexity of HHL, this error is assumed to be solely from inadequacy in the number of clock register qubits [31]. In the above expression, we have explicitly shown the dependence of condition number, sparsity, and error on $\mathcal{N}$ to stress its importance in the context of our study. The scaling in $\kappa(\mathcal{N})$ and $\epsilon(\mathcal{N})$ are usually considered as typical drawbacks, and variants of HHL and subsequent QLSs improve on one or both of these aspects.
@@ II.3 Variants of HHL and other QLSs
We now list four variants of HHL that we consider for this work, all of which focus on reducing $\kappa$ scaling:
  • HHL with Amplitude Amplification (HHL-AA) [31, 12]: This is often assumed when one discusses HHL, but since the subroutine adds significant depth to the HHL quantum circuit (for example, see Ref. [46]), we keep it distinct from the original HHL circuit. The benefit that the variant offers is a reduction of complexity in $\kappa(\mathcal{N})$, from $\kappa(\mathcal{N})^{3}$ to $\kappa(\mathcal{N})^{2}$. Practically, this has the effect of reducing the number of shots in an HHL calculation.
  • HHL with Variable Time Amplitude Amplification (HHL-VTAA) [2]: This approach improves over HHL-AA and can be thought of as its generalization. The method reduces the complexity further to $\kappa(\mathcal{N})\mathrm{log}^{3}(\kappa(\mathcal{N}))$, but trades off in $\epsilon(\mathcal{N})$ scaling (see Table 1).
  • Psi-HHL [64]: This recently introduced approach reduces the complexity in $\kappa$ to its optimal scaling, that is, $\kappa(\mathcal{N})$, and with little increase in circuit depth, for cases with large $\kappa$ values.
@@
[…]
We now list the other QLSs that we consider besides HHL and its variants:
  • Phase randomisation method [62]: The method is inspired by the adiabatic quantum computing model, and employs evolution randomisation. The method scales as
  • AQC(exp) method [3]: The work demonstrates solving system of linear equations within the adiabatic quantum computing framework. This method too scales as
@@
    and thus is near-optimal in scaling of both condition number and precision, but unlike the CKS algorithm, the use of the expensive VTAA step is avoided. However, the minimum gap between the ground state and excited state of the time-dependent Hamiltonian that is varied adiabatically plays a strong role in determining the runtime of the algorithm [3]. Thus, only problems where the spectral gap shrinks polynomially in system size retain the speedup offered by this QLS.
@@
    The dream QLS serves as a benchmark to how much of an advantage we can get in the best case scenario, and thus subsumes all other QLSs that we do not consider. We assume in defining this solver that one cannot go below $\kappa(\mathcal{N})$ [31] , $\sqrt{s(\mathcal{N})}$ [47], and $\mathrm{log}(1/\epsilon(\mathcal{N}))$ [46] in its complexity expression.
@@
HHL [31] $\mathcal{O}\textbf{\bigg(}\mathrm{log}(\mathcal{N})s(\mathcal{N})^{2}\kappa(\mathcal{N})^{3}\frac{1}{\epsilon(\mathcal{N})}\textbf{\bigg)}$
HHL-AA [31, 12] $\mathcal{O}\textbf{\bigg(}\mathrm{log}(\mathcal{N})s(\mathcal{N})^{2}\kappa(\mathcal{N})^{2}\frac{1}{\epsilon(\mathcal{N})}\textbf{\bigg)}$
HHL-VTAA [2] $\mathcal{O}\textbf{\bigg(}\mathrm{log}(\mathcal{N})s(\mathcal{N})^{2}\kappa(\mathcal{N})\mathrm{log}^{3}\bigg(\frac{\kappa(\mathcal{N})}{\epsilon(\mathcal{N})}\bigg)\frac{1}{\epsilon(\mathcal{N})^{3}}\mathrm{log}^{2}\bigg(\frac{1}{\epsilon(\mathcal{N})}\bigg)\textbf{\bigg)}$
Psi-HHL [64] $\mathcal{O}\textbf{\bigg(}\mathrm{log}(\mathcal{N})s(\mathcal{N})^{2}\kappa(\mathcal{N})\frac{1}{\epsilon(\mathcal{N})}\textbf{\bigg)}$
Phase randomisation method [62] $\mathcal{O}\textbf{\bigg(}\mathrm{log}(\mathcal{N})s(\mathcal{N})\kappa(\mathcal{N})\mathrm{log}(\kappa(\mathcal{N}))\frac{1}{\epsilon(\mathcal{N})}\textbf{\bigg)}$
CKS algorithm [15] $\mathcal{O}\textbf{\bigg(}\mathrm{log}(\mathcal{N})s(\mathcal{N})\kappa(\mathcal{N})\mathrm{polylog}\bigg(s(\mathcal{N})\kappa(\mathcal{N})\frac{1}{\epsilon(\mathcal{N})}\bigg)\textbf{\bigg)}$
AQC(exp) method [3] $\mathcal{O}\textbf{\bigg(}\mathrm{log}(\mathcal{N})s(\mathcal{N})\kappa(\mathcal{N})\mathrm{polylog}\bigg(s(\mathcal{N})\kappa(\mathcal{N})\frac{1}{\epsilon(\mathcal{N})}\bigg)\textbf{\bigg)}$
@@
Table 1 presents the expression for the runtime complexities of each of the aforementioned quantum algorithms in its second column. We assume throughout hereafter that $\epsilon^{-1}\sim\mathrm{log}(\mathcal{N})$, so that the desired polylog runtime complexity of a QLS is retained. We also drop the $\mathcal{O}$ hereafter when we discuss complexity expressions, for brevity.
@@
The specific functional forms for ${f}(s(\mathcal{N}))$ and ${g}(\kappa(\mathcal{N}))$ depend on the QLS chosen. On the other hand, $\kappa(\mathcal{N})$ and $s(\mathcal{N})$ depend on the choice of application, and they determine the degree of quantum advantage that one obtains for an application.
We now make some general observations. When $R(\mathcal{N})>1$, a QLS under consideration performs better than CLS. That is, when
@@
That is, if an $(\mathcal{N}\times\mathcal{N})$ matrix $A$ is ill-conditioned enough and/or is sufficiently dense such that the product $\kappa(\mathcal{N})^{5/2}s(\mathcal{N})$ is equal to or greater than the right hand side, then we cannot expect an advantage from HHL. If we go to an improved variant such as Psi-HHL, the product $\kappa(\mathcal{N})^{1/2}s(\mathcal{N})$ can be less than the right hand side for a matrix $A$ that is relatively denser and/or relatively more ill-conditioned. The condition for the phase randomisation method is particularly interesting, as its sparsity scaling is the same as that of CLS, and thus it is only the condition number scaling that matters when we compare the two approaches. Similarly, for the dream QLS case, as long as $\kappa$ and $s$ were to scale the same way, independent of their behaviours with $\mathcal{N}$, the runtime ratio is $\mathcal{N}/\mathrm{log}(\mathcal{N})$ thus resulting in an exponential separation in the runtimes. We also note that for the cases of the CKS and the AQC(exp) algorithms, it is not possible to separate out $\kappa(\mathcal{N})$ and $s(\mathcal{N})$ to one side due to the functional form for their complexities. Lastly, it is worth adding that since the complexity expressions contain a polylog in them, we will, for the purposes of our numerical analyses in the subsequent sections, consider three cases for the CKS and AQC(exp) algorithms for polylog: log, log2, and log3 behaviours. In the subsequent figures as well as main text, we use a shorthand notation: for example, the AQC(exp) algorithm with the polylog chosen to be log is denoted as AQC(1), and so on.
We now list the expressions for $R(\mathcal{N})$ for HHL, CKS(1)/AQC(1), and the dream QLS:
  1. 1.
    HHL:
    $R(\mathcal{N})_{1}=\frac{\mathcal{N}}{\mathrm{log}(\mathcal{N})}\frac{\mathrm{log(log(\mathcal{N})})}{\mathrm{log(}\mathcal{N})\kappa^{5/2}s},$ (1)
  2. 2.
    : CKS(1)/AQC(1):
    $R(\mathcal{N})_{2}=\frac{\mathcal{N}}{\mathrm{log}(\mathcal{N})}\frac{\mathrm{log(log(}\mathcal{N}))}{\sqrt{\kappa}\mathrm{log(log(}\mathcal{N})s\kappa)},$ (2)
    and
  3. 3.
    Dream QLS:
    $R(\mathcal{N})_{3}=\frac{\mathcal{N}}{\mathrm{log}(\mathcal{N})}\frac{\sqrt{s}}{\sqrt{\kappa}}.$ (3)
We remark here that $\frac{\mathcal{N}}{\mathrm{log}(\mathcal{N})}$ is an exponential separation in $\mathrm{log}(\mathcal{N})$. To retain the exponential behaviour in $R(\mathcal{N})_{1}$ under the assumption that $\epsilon^{-1}\sim\mathrm{log}(\mathcal{N})$, the term $\kappa^{5/2}s$ must strictly grow less than $\mathcal{N}$. With $R(\mathcal{N})_{2}$, it is sufficient for $\kappa$ to grow less than $\mathcal{N}^{2}$. However, with $R(\mathcal{N})_{3}$, to observe exponential advantage, the growth of $\kappa/s$ should be less than $\mathcal{N}^{2}$. We also note at this juncture that our analysis is purely based on graph theoretic considerations and neglects implementation overheads such as the number of physical qubits given a quantum error correcting code, cost of magic state distillation, etc. In view of the possibility of such overheads overriding advantages less than quadratic (for example, see [5]), we consider, for example, logarithmic speedup, as no advantage in this study.